# LLMLight: Large Language Models as Traffic Signal Control Agents

### v4

@@ Appendix A Appendix
A.1. Visualizations of Intersection, Lanes, and Signal Phases
@@ A.2.1. Transportation Methods
  • Maxpressure (Varaiya, 2013): The state-of-the-art heretics traffic signal control method in the transportation field, which greedily relieves vehicles on the lanes with the highest pressure (a variable derived from the difference between upstream and downstream queue lengths).
@@ A.2.3. Large Language Models
  • Llama 2 (Touvron et al., 2023): This is a series of pre-trained and finely-tuned large language models (LLMs) developed by the AI group at Meta, ranging in scale from 7 billion to 70 billion parameters. They are adaptable for various natural language generation tasks. In our experiments, we use Llama-2-13B-chat-hf and Llama-2-70B-chat-hf in our experiments.
@@
  • GPT-4 (OpenAI, 2023a): As the newest advancement in LLM by OpenAI, this SOTA foundation model is more reliable and creative compared to GPT-3.5. It excels in handling significantly more nuanced instructions and tasks. We utilize gpt-4-0613 API in our experiments.
@@
A.3. Failure of Other LLMs
To delve into the failure of other LLMs, we present the reasoning processes of Llama-2 and ChatGPT-3.5 in Figure 10. Table 3 shows the detailed traffic conditions. This scenario shows heavy congestion in the north and south sections, characterized by a notable accumulation of queuing vehicles in the left-turn lane and a considerable influx of approaching vehicles in the go-through lane. The Llama-2 models exhibit a common limitation in traffic management expertise and instruction-following (they disregard the provided commonsense knowledge). Although they understand the optimal traffic signal needs to address the most congested lanes of the intersection, Llama-2 opts to prioritize approaching vehicles, neglecting the more critical factors, queuing vehicles. On the other hand, ChatGPT-3.5 manifests a substantial logical flaw by asserting that releasing lanes with the lowest vehicle count is most effective in enhancing traffic conditions. This observation may suggest a pronounced hallucination issue (Ji et al., 2023; Huang et al., 2023) of ChatGPT-3.5 in the context of traffic signal control.
A.4. Visualizations of Road Networks

### v5

@@ Appendix A Appendix
A.1. Settings of Intersection, Lanes, and Signal Phases
@@ A.2.1. Transportation Methods
  • Maxpressure (Varaiya, 2013): The state-of-the-art transportation-based method, which greedily relieves vehicles on the lanes with the highest pressure (a variable derived from the difference between upstream and downstream queue lengths).
@@ A.2.3. Large Language Models
  • Llama (Touvron et al., 2023): This is a series of pre-trained and finely-tuned large language models (LLMs) developed by Meta, ranging in scale from 7B to 70B parameters. They are adaptable for various tasks related to natural language generation. We use the second- and third-generation, Llama-2 and Llama-3, in our experiments.
  • Qwen2 (Bai et al., 2023): Alibaba’s Qwen2 series are LLMs with parameters ranging from 0.5B to 72B, supporting up to 128K tokens context, enhancing capabilities in natural language understanding, code writing, and multi-language processing.
@@
  • GPT-4 (OpenAI, 2023a): The newest advancement in LLM by OpenAI. It is more reliable and creative compared to GPT-3.5. It excels in handling significantly more nuanced instructions and tasks. We utilize gpt-4-0613 API in our experiments.
@@
Figure 9. Prototype system.
A.3. Prototype System
Figure 9 shows a snapshot of our demo system. We provide users with real-time visualizations of the current traffic conditions on the right, along with statistical data on vehicles in the upper left corner. Additionally, the system further presents the reasoning behind LightGPT’s decisions through a message box in the interface’s lower left section.
Table 4. The annual deployment cost (in USD) of backbone LLMs. LightGPT-8B and LightGPT-13B states LightGPT (Llama3-8B) and LightGPT (Llama2-13B), respectively.
City scale Small Medium Large Metropolitan
# of traffic lights 100 500 1000 5000
Model GPT-4 LightGPT-13B LightGPT-8B GPT-4 LightGPT-13B LightGPT-8B GPT-4 LightGPT-13B LightGPT-8B GPT-4 LightGPT-13B LightGPT-8B
# of machines - 2 1 - 10 5 - 20 10 - 100 50
Annual cost \$1,680K \$44.91K \$22.45K \$8,400K \$224.54K \$112.27K \$16,800K \$449.08K \$224.54K \$84,100K \$2,250K \$1,120K
Table 5. The training budget (in USD) of LightGPT.
Data Collection IFT (1 GPU) CGPR (4 GPUs) Sum
Cost \$135.53 \$11.60 \$5.52 \$152.65
Table 6. The efficiency (seconds per batch) of backbone LLMs.
# of traffic intersections (batch size) 1 5 10 15 20
GPT-4 8.60s 9.23s 11.69s 11.35s 12.95s
LightGPT (Llama2-13B) 3.53s 4.61s 5.31s 5.74s 6.62s
LightGPT (Llama3-8B) 2.52s 3.30s 3.85s 4.15s 4.69s
A.4. Cost-effectiveness of LightGPT
To assess the cost-effectiveness of our proposed LightGPT with the most widely used and advanced LLM, GPT-4. We report the annual deployment cost in different city scales, budget in data collection and model training, and efficiency evaluation of backbone LLMs in Tables 4, 5, and 6, respectively. These calculations were performed on a cloud computing machine with 4 × A800-80GB GPUs, 56 cores, 400GB of memory, and a 1TB SSD. Please note that our model weights have been open-sourced. With its exceptional generalization ability, it can be directly deployed in most unseen traffic environments without further fine-tuning.
For real-time efficiency in deployment, we set the number of parallelly controlled intersections to 20 for LightGPT (Llama3-8B) and 10 for LightGPT (Llama2-13B). When controlling asynchronously, one machine with LightGPT (Llama3-8B) can serve up to 100 intersections, while LightGPT (Llama2-13B) can serve up to 50. We can observe that LightGPT is the most cost-effective solution, as it achieves the best performance in both effectiveness (see Section 4) and real-time traffic control, while also having the lowest deployment cost. Further efficiency and cost-effectiveness improvements are planned for future work.
A.5. 24-hour Cycle Experiment
We extend our experiment to a full 24-hour period to examine model performance under daily long-term traffic fluctuations, including distinct morning and evening rush hours. Detailed traffic flow distribution of this dataset can be found in Figure 11. All models are pre-trained using the Jinan 1 dataset. The results are presented in Figure 10. Consistent with previous experiments, we observe that the performance of RL methods drops significantly when traffic flow shifts across different times of the day, as evidenced by their poorer results compared to Maxpressure. As concluded in Subsection 4.7.2, RL methods result in significantly longer waiting times than Maxpressure and LLMLight, due to their focus on minimizing overall queuing vehicles, which may possibly extend the waiting times of queues with fewer vehicles. In contrast, our proposed method consistently ensures both the shortest travel and waiting times, demonstrating its robustness and real-world applicability in diverse and fluctuating traffic conditions. Moreover, LLMLight significantly outperforms the widely used FixedTime method in real-world scenarios, further underscoring the practical benefits of implementing our approach in real-world deployments.
Figure 10. The performance in a 24-hour period.
Figure 11. The traffic flow distribution in the 24-hour.
A.6. Human Expert Evaluation
We conduct a human evaluation across five randomly selected traffic scenarios to evaluate LightGPT’s reasoning and decision-making capabilities under diverse traffic environments. Feedback was solicited from a diverse group: four traffic police officers from the Kunming Southeast Ring Expressway Jinsong Traffic Police Brigade in Yunnan and the Zhongshan Traffic Police Detachment in Guangdong, China, nine experienced drivers, and two AI specialists. The evaluation results, the studied traffic scenarios, and the reasoning process of LightGPT are presented in Figure 12-16. The consensus from these evaluations was overwhelmingly positive, with participants praising its applicability and notable interpretability in real-world contexts.
A.7. Other Deployment Considerations
A.7.1. Challenges in Real-world Deployment
We outline several challenges that may arise during real-world deployment: (1) Integration with existing systems: LLMLight relies on traffic cameras or ground sensing coils for data collection. While these devices are common in most traffic scenarios, modifications might be needed to integrate our framework seamlessly with current systems. (2) Regulatory compliance: Our experiments have primarily been conducted in China and the U.S., which have different regulatory frameworks compared to other countries. Variations in regulations, such as driver seat positioning and road rules, may require additional adjustments to tailor our model for deployment in various regions. (3) Impact on driver behavior: Unlike traditional traffic signal control (TSC) systems, which follow a predetermined sequence, intelligent TSC systems like LLMLight introduce unpredictability in signal switching. This unpredictability could potentially influence driving behaviors. Further adjustments to address these challenges are planned for future work.
A.7.2. Data Privacy
For real-world deployment, we plan to utilize existing infrastructure like traffic cameras and ground sensing coils for observation feature collection. Our system solely extract non-personal data (i.e., vehicle counts) to obtain traffic conditions. Any images captured will be securely stored on the traffic department’s server and will not be made public, ensuring the maintenance of privacy integrity.
@@ A.7.2. Data Privacy
Table 8. The performances of LLMLight with the original LLMs (without fine-tuning) and LightGPTs on the Jinan 1 and Hangzhou 1 datasets.
Methods Jinan 1 Hangzhou 1
ATT AQL AWT ATT AQL AWT
LLMLight (with Generalist LLMs)
Qwen2-0.5B 1306.46 1771.49 953.10 1124.84 789.44 880.79
Qwen2-7B 618.86 790.53 469.08 788.98 497.46 616.32
Llama2-7B 1115.97 1580.44 896.20 973.50 664.58 819.00
Llama3-8B 983.48 1236.04 830.66 1136.17 643.44 806.26
Llama2-13B 508.72 567.61 169.67 722.72 418.64 311.55
LLMLight (with LightGPTs)
LightGPT (Qwen2-0.5B) 296.71 193.93 46.80 328.11 70.90 84.49
LightGPT (Qwen2-7B) 275.92 163.34 48.56 313.37 59.03 50.65
LightGPT (Llama2-7B) 275.11 161.35 46.38 314.24 59.59 39.66
LightGPT (Llama3-8B) 275.10 161.92 48.25 311.72 58.40 47.60
LightGPT (Llama2-13B) 274.03 159.39 43.24 310.78 56.93 38.64
A.8. Failure of Other LLMs
To explore the shortcomings of other LLMs, we analyze the reasoning processes of Llama-2 and ChatGPT-3.5, as illustrated in Figure 17. Table 7 details the traffic conditions, highlighting severe congestion in the north and south sections, with notable queues in the left-turn lane and a heavy influx in the go-through lane. Llama-2 models commonly show limitations in traffic management expertise and instruction-following, often disregarding provided commonsense knowledge. Although they recognize the need to address the most congested lanes, Llama-2 prioritizes approaching vehicles, overlooking the critical queuing vehicles. Conversely, ChatGPT-3.5 exhibits a major logical flaw, claiming that releasing lanes with the lowest vehicle count is most effective for improving traffic conditions. This observation suggests a significant hallucination issue (Ji et al., 2023; Huang et al., 2023) in ChatGPT-3.5’s traffic signal control reasoning. The performances of more generalist LLMs are detailed in Table 8.
A.9. Experiment on Policy Network Learning by Using the Actor-Critic Paradigm
We explore using the actor-critic paradigm, specifically Proximal Policy Optimization (PPO) (Schulman et al., 2017), to learn the traffic control policy. Similar to reinforcement learning with human feedback (RLHF) (Christiano et al., 2017), we integrate a multi-layer perceptron (MLP) into the final decoder layer of the LLM as the value function head. The reward function is based on the negative queue length at the target intersection. PPO is employed to facilitate automatic exploration and learning of the optimal control policy and value function. However, this approach struggled to converge and yielded unsatisfactory results, as reported by (Wang et al., 2023a). Consequently, we adopted our proposed critic-guided policy refinement (CGPR), which uses the critic model for supervised alignment training, ensuring robust performance. Table 9 compares the PPO algorithm and our proposed CGPR on the Jinan 1 and Hangzhou 1 datasets.
Table 9. The comparison between PPO and CGPR.
Methods Jinan 1 Hangzhou 1
ATT AQL AWT ATT AQL AWT
LLMLight (LightGPT (PPO)) 275.35 164.64 57.07 315.32 60.63 50.63
LLMLight (LightGPT (CGPR)) 274.03 159.39 43.24 310.78 56.93 38.64
Table 10. The comparison between the SOTA RL and LLMLight without agent cooperation.
Methods Hangzhou 1 Hangzhou 2
ATT AQL AWT ATT AQL AWT
Advanced-CoLight (W/O COOP) 310.48 58.02 54.54 332.08 192.97 88.51
LLMLight (LightGPT (Llama2-13B)) 310.78 56.93 38.64 330.71 189.09 64.16
A.10. Comparison with the SOTA RL-based Method Under a Fair Setting
Despite Advanced-CoLight’s marginal advantage in the Hangzhou datasets, note it requires additional communication among neighboring intersections. However, LLMLight achieves competitive results by leveraging observation features solely from the target intersection. We conducted an ablation study, removing the multi-agent cooperation component from Advanced-CoLight, denoted as Advanced-CoLight (W/O COOP). The experiment results are reported in Table 10. In this comparison, LLMLight showcased superior performance in ATT, AQL, and AWT on the Hangzhou dataset. In summary, LLMLight achieves better performance with the same input, and we believe LLMLight can be further improved in the future by incorporating multi-agent cooperation mechanisms.
A.11. Visualizations of Road Networks
@@ A.12. Prompt Template
Figure 12. Human expert evaluation: Case 1.
Figure 13. Human expert evaluation: Case 2.
Figure 14. Human expert evaluation: Case 3.
Figure 15. Human expert evaluation: Case 4.
Figure 16. Human expert evaluation: Case 5.